The following data are for the two products produced by Tadros Company. Product A Product B Direct materials $15 per unit $241. Find the Activity Rate.

2. Find the Assigning Overhead Cost.
Answer #1

Solution:

1. Calculation of Activity Rate.

Engineering support = Overhead cost / Engineering Modifications

Engineering support = $24,500 / (12+58) = $24,500 / 70

Engineering support = $350 per Engineering Modifications

Electricity = Overhead cost / Machine Hours

Machine Hours for Product A = Machine Hours per unit \times Number of Units

Machine Hours for Product A = 0.1 Machine Hours per unit \times 10,000 Units = 1,000 Machine Hours

Machine Hours for Product B = Machine Hours per unit \times Number of Units

Machine Hours for Product B = 1.2 Machine Hours per unit \times 2,000 = 2,400 Machine Hours

Electricity = Overhead cost / Machine Hours = $34,000 / (1,000 + 2,400) = $34,000 / 3,400 = 10 per Machine Hour.

Electricity = 10 per Machine Hour.

Set up cost = Overhead cost / Number of Batches

Set up cost = $52,500 / (125 + 225) = $52,500 / 350

Set up cost = $150 per Batch

Activity rate for Non manufacturing cost

Customer service = Overhead cost / Number of customers

Customer service = $81,000 / (500 + 400) = $81,000 / 900

Customer service = $90 per customer

2. Calculation of Assigning Over head cost

Particulars Rate Product A Product B
Engineering Support $350 $350 \times 12 = $4,200 $350 \times 58 = $20,300
Electricity $10 $10 \times 1,000 = $10,000 $10 \times 2,400 = $24,000
Set up cost $150 $150 \times 125 = $18,750 $150 \times 225 = $33,750
Total Manufacturing over head cost $32,950 $78,050

Assigning non manufacturing overhead cost

For product A = Number of customers \times rate per customer = 500 \times $90 = $45,000

For product B = Number of customers \times rate per customer = 400 \times $90 = $36,000